OpenStudy (anonymous):

Need quick help on half this problem? Bailey pulls a letter tile from a bag that contains one of each letter of the alphabet. Match each description of the tile with its probability.

OpenStudy (anonymous):

1. a vowel in her name 2. a letter in her name or V 3. a letter in her name 4. a letter in her name or a consonant (Assume vowels are A, E, I, O, and U.) a. 12/13 b. 3/13 c. 7/26 d. 3/26 ---------------------------------------- I have matched "a vowel in her name" with 3/26 also matched "a letter in her name or v" with 7/26 however, I would believe that a letter in her name would be 6/26 and the last one would be 24/26

OpenStudy (debbieg):

Hmmm..... maybe 6/26 is equivalent to one of the fractions given as options? Maybe 24/26 also? ;)

OpenStudy (anonymous):

6/26 is equivalent to 12/13 and 24/26 is equivalent to 3/13

OpenStudy (anonymous):

i suppose that would actually work

OpenStudy (debbieg):

noooo..... reduce 6/26. What's the common factor?

OpenStudy (anonymous):

oh sorry i've been doing math all friggen day. im burnt out. it reduces to 3/13

OpenStudy (debbieg):

remember, for two fractions to be "equivalent", you must be able to multiply the num'r and den'r of one of them by the SAME NUMBER, to produce the other fraction. E.g.: \[\Large \dfrac{ 21 }{ 28 }=\dfrac{ 7\cdot 3 }{ 7 \cdot4 }=\dfrac{ 3 }{ 4 }\]

OpenStudy (anonymous):

yes

OpenStudy (debbieg):

ok, right. You had them switched. :) And yes, that WILL work - it's the correct answer because fractions should be fully reducced.

OpenStudy (anonymous):

oh yeah i forgot about reducing them. thank you very much for your time and help